OpenStudy (anonymous):

∫x3/(x2+2x+1)=∫x−2+(3x+2)/(x+2)2dx Would my next step of partial fraction evaluation be correct?

OpenStudy (anonymous):

\[\int\limits_{}^{} x^{3} / (x^{2} + 2 x + 1) = \int\limits_{} x -2 + (3x + 2 ) / (x+2)^2 dx\] Would my next step of partial fraction evaluation be correct?

OpenStudy (anonymous):

The denominator should not be \[(x+2)^2\], though it can be factored.

OpenStudy (anonymous):

the denominator should be x^2 + 2x +1 ?

OpenStudy (anonymous):

You can write that or \[(x+1)^2\]

OpenStudy (anonymous):

thank you